# Fine-Grained Vision-Language Modeling for Multimodal Training Assistants in Augmented Reality

### v1

@@ 5.3 Case Study
GPT-4o Ensure the transparent mailbox front 2x2 is securely attached by inserting the small shaft into the round brick 1x1 before proceeding with the next step.
Table 3: An example of scene understanding results. Words highlighted in orange represent manipulation verbs, red and blue indicate noun entities and their adjectives, and purple denotes positions. The word with a background color orange indicates the task-specific token and the words with a background color green is obtained from the previous step’s textual instruction manual.
@@
Table 3 present an example showcasing the intuitive results of the prevailing VLMs on scene understanding. We obtain the following observations: First, most models are able to correctly identify the core theme entity “mailbox front” (highlighted in green), with the exception of Otter and GPT-4o when performing scene understanding. However, generating a complete set of properties (e.g., “transparent” highlighted in blue and size “2x2” in pink) to describe these entities accurately and without ambiguity remains a significant challenge. Second, providing the correct assembly operations with precise positional instructions is highly challenging. Although LlaVA and MiniGPT-v2 generate informative instructions, they are not factually accurate in terms of the assembly process.
6 Related Work
6.1 Vision-Language Datasets in AR
Dataset COIN HowTo100M TEACh MIMIC-IT HoloAssist EPIC-KITCHENS Assembly101 Cross-task RareAct LEGO-VLM (Ours)
T1
T2
T3
Size 11,827 23,611 3,215 2.8M 350 89,977 4,321 4,713 7,607 35,612
Table 4: Comparison of datasets for assembly tasks towards AR training regarding the capabilities of tasks T1, T2, and T3. The symbols ✓and ✗indicate the presence or absence of each capability.
@@ 6.1 Vision-Language Datasets in AR
Various datasets have been developed to enhance AR assembly tasks by supporting key capabilities T1, T2, and T3. A comparative summary is presented in Table 4.
@@
COIN provides a strong foundation for sequential task analysis with comprehensive annotations, aiding research in multimodal learning (Tang et al., 2019). HoloAssist captures egocentric human-AI interactions using mixed-reality headsets, offering valuable real-world insights (Wang et al., 2023b). HowTo100M, with its extensive collection of 136 million video clips and transcribed narrations, is highly effective for text-to-video retrieval but lacks fine-grained object annotations (Miech et al., 2019). TEACh focuses on interactive dialogues in domestic environments, improving dialogue modeling but not prioritizing detailed object detection (Padmakumar et al., 2021).
@@
While some datasets excel in action recognition, they often lack the procedural depth necessary for effective training. Assembly101, with over 4,000 toy assembly videos, does not include step-by-step instructional details (Sener et al., 2022). CrossTask facilitates weakly supervised learning by leveraging narrations and step lists but lacks temporal annotations to clarify action sequences (Zhukov et al., 2019). EPIC-KITCHENS provides extensive annotations on unscripted kitchen activities, yet it does not offer structured procedural guidance (Damen et al., 2018). RareAct presents unique interactions, challenging models to interpret complex actions without explicit instructions (Miech et al., 2020).
@@
Stanescu et al. (2023a) introduce a state-aware prior that significantly improves object detection in assembly tasks like furniture and Lego construction. However, it lacks textual descriptions of object states and does not address fine-grained detection (Stanescu et al., 2023b). Meanwhile, MIMIC-IT contributes 2.8 million multimodal instruction-response pairs, enriching conversational modeling but lacking scenario-specific descriptions that would enhance task guidance (Li et al., 2023a).
@@
6.2 Towards Fine-grained VLMs
@@
Recent advancements in vision-language models (VLMs) have broadened their applications, enhancing accessibility, search indexing, and interactive content retrieval (Fan et al., 2024; Zhang et al., 2024; Du et al., 2022; Gan et al., 2022). These models generate captions, describe images (Zhou et al., 2020; Hu et al., 2022), facilitate visual QA (Bazi et al., 2023). For instance, they can accurately identify objects in images based on phrases like “the red car” (Subramanian et al., 2022). Furthermore, these models facilitate conversational interactions about visual elements (Chen et al., 2022), which makes virtual assistants more adept at handling inquiries related to images and videos, which makes virtual assistants more adept at handling inquiries related to images and videos. Additionally, VLMs can support multilingual content (Gwinnup & Duh, 2023), benefiting assistive technologies and education (Chi et al., 2020; Wang et al., 2024a).
@@
Despite progress, VLMs often lack fine-grained modeling for scene understanding, object recognition, and error detection. While CLIP demonstrates strong generalization (Radford et al., 2021), models like EfficientVLM (Wang et al., 2023a), MiniGPT-v2 (Chen et al., 2023), and Qwen-VL (Bai et al., 2023) struggle with intricate tasks. OSCAR (Li et al., 2020) and VisionLLM (Wang et al., 2024b) improve image-text alignment but face challenges in context-specific adaptation. The Otter model advances sequential task management yet falls short in detailed analysis (Li et al., 2023b).
@@
Efforts in procedural video representation (Zhong et al., 2023) and instructional task graphs (Ashutosh et al., 2024) highlight potential VLM integration but reveal gaps in error correction. To address these challenges, we propose a vision-language architecture optimized for fine-grained tasks, advancing training and benchmarking of VLMs.
7 Conclusion
@@
Vision-language models (VLMs) have significantly advanced in recent years, enabling AI systems to interpret and generate textual descriptions of visual content. However, despite their success in general vision-language tasks, these models often struggle with fine-grained understanding, particularly in structured instructional scenarios. This research explores fine-grained vision-language modeling for manual-guide AR training, focusing on scene understanding, object detection, and state detection—areas where VLMs often struggle. We developed a specialized dataset from LEGO instruction manuals, designing fine-grained tasks to evaluate VLM performance in tracking assembly sequences and interpreting instructions. Through a one-for-all architecture, we assessed existing models and found them lacking in precision for instructional tasks, highlighting the need for improvement. Future work will analyze model failures and develop enhanced VLMs with better accuracy and contextual understanding. Beyond technical advancements, our research aims to empower blind and visually impaired individuals by enabling AI-driven learning tools, promoting greater accessibility and independence.
Appendix A Analysis
A.1 Impact of Query Construction
We explore the impact of query construction on the QWEN-VL model in the early phase to identify the optimal configuration for achieving the best inference results.
Setting T1 T2 T3
Instruction Example T_ACC $\uparrow$ ROUGE $\uparrow$ BLEU $\uparrow$ E_ACC $\uparrow$ IOU $\uparrow$ F1-State $\uparrow$ FPR $\downarrow$
11.90 29.57 5.19 32.41 25.60 39.77 97.99
6.84 24.63 4.58 5.56 22.01 34.00 97.50
10.43 23.00 4.24 26.37 15.55 34.50 97.20
5.27 11.02 1.30 3.70 0.10 25.50 98.40
Table 5: A study of query construction results across various tasks and metrics. The symbol $\uparrow$ indicates higher is better, $\downarrow$ indicates lower is better, and values in bold represent the highest results performance.
To this end, we conducted experiments on various query settings to determine which components most significantly influence performance. The model was chosen for this study due to its strong results under our base setting, which included both instruction directives and in-context examples.
As shown in Table 5, our findings highlight the critical role of query design across different tasks. The results indicate that providing both instructions and examples yields the best overall performance, demonstrating their complementary roles in guiding inference.
In the scene understanding task (T1), the presence of instructions plays a critical role in improving accuracy. When instructions are included, theme accuracy (T_ACC) increases significantly, suggesting that clear guidance helps the model better understand scene context and thematic elements. Furthermore, both ROUGE and BLEU scores, which evaluate the structure and coherence of generated outputs, improve when at least one of the two components (instructions or examples) is present. The lowest performance in this task occurs when neither component is provided, confirming that structured input is essential for meaningful inference.
In the object detection task (T2), the results suggest that examples contribute more to entity recognition and output formatting than instructions alone. The entity accuracy (E_ACC) drops sharply from 32.41% to 5.56% when examples are removed, even when instructions remain. Similarly, IoU, which measures how well the output adheres to the expected format, significantly decreases in the absence of examples. This highlights that examples provide concrete reference points, enabling the model to generate well-structured and accurate entity labels.
In the state detection task (T3), despite improvements in other metrics with enhanced query settings, the false positive rate (FPR) remains consistently high (above 97%) across all conditions. This suggests that the model has a strong affirmative bias, likely defaulting to generating positive responses regardless of query refinements. While instructions and examples improve structured outputs, they do not sufficiently mitigate the issue of excessive false positives.
These findings emphasize the importance of query design in optimizing model performance. Combining instructions with instruction directives and in-context examples is crucial for achieving high accuracy, structured outputs, and reliable entity recognition. However, the persistent issue of high FPR suggests the need for further refinements, such as introducing negative examples, confidence calibration, or additional fine-tuning strategies to reduce overconfidence in affirmative responses. Future work should explore these approaches to enhance the model’s robustness and reliability in diverse inference tasks.
A.2 Case Study of Object Detection
Text Image
Query [object] Please identify and highlight the positions of 1 light purple brick 2x6. Providing the positions in the format: <p>object</p> <Xleft><Ytop><Xright><Ybottom>, with X and Y coordinates normalized to [0,100]. <Xleft> and <Ytop> for the top-left corner. <Xright> and <Ybottom> for the bottom-right corner. [Uncaptioned image]
Reference <p>1 light purple brick</p> <58><15><92><35>
mPLUG-OWL2 <p>1 light purple brick</p> <125><48><45><97>
Fine-tuned <p>1 light purple brick</p> <60><24><69><35>
BLIP2 <p>1 light purple brick</p> <57><15><82><35><p>and<p>1 dark purple brick 2x6</p> <82><25><93><36>
Fine-tuned <p>1 light purple brick</p> <58><15><92><35>
LLaVA The light purple brick is located at the center of the image, surrounded by yellow and black bricks.
Fine-tuned <p>1 light purple brick</p> <62><15><87><35>
QWEN-VL <20><61><30><78>
Fine-tuned <p>1 light purple brick</p> <57><15><82><35>
InstructiBLIP <p>1 light purple brick 2x6</p> <57><15><82><35>
Fine-tuned <p>1 light purple brick</p> <58><15><92><35>
MiniGPT-v2 <p>1 light purple brick</p> <62><20><88><40>
Fine-tuned <p>1 light purple brick</p> <57><17><85><37>
Otter The 1 light purple brick 2x6 is on the left side of the image.
MiniGPT-4 Sure! Here are the positions of the light purple brick in the image: Xleft = 0, Ytop = 25, Xright = 30, Ybottom = 50.
GPT-4o from PIL import Image # Load the image image_path $\dots$
Table 6: An example of object detection results. The object is highlighted in light purple, and its position is highlighted in blue.
Table 6 present an example of object detection task. We obtain the following observations: First, only InstructiBLIP and Otter generate the full description of the object “1 light purple brick 2x6” without disambiguation. Second, none of the models accurately predict the precise positions as outlined in the reference.
Appendix B Enlightening Discussion
B.1 Limitations
The quality of object and state detection in our dataset heavily depends on the performance of current VLMs. While we have assessed the data quality, it is important to note that this is a synthetic dataset with only limited human intervention for corrections. In the object detection task, the model generates coordinates used to highlight detected objects with bounding boxes, though these coordinates are not meant for direct user interaction. In contrast, the outputs of the state detection and scene understanding tasks are text-based and designed to be user-friendly. However, we have not yet accounted for 3D scenes or objects. We evaluated a few existing VLMs as benchmarks, but their performance remains suboptimal. Future work will focus on improving fine-grained VLMs to enhance their effectiveness in these tasks.
B.2 Broader Impact
Our work has significant implications for accessibility and inclusive learning, particularly in assisting blind and visually impaired users (Gadiraju et al., 2020; Hao et al., 2024; Karamolegkou et al., 2024; Zhao et al., 2024). By integrating VLMs into conversational training assistants, we aim to create equitable learning opportunities in AR-based training environments. This could enable individuals with visual impairments to engage in hands-on assembly tasks through AI-powered guidance, reducing barriers to education and skill development.
Beyond accessibility, our research contributes to industrial applications, particularly in manufacturing, robotics, and assembly training. Fine-grained vision-language reasoning can enhance AR-based instruction systems, reducing errors, improving efficiency, and enabling more intuitive human-AI collaboration in high-stakes environments like car manufacturing and industrial product assembly. By bridging the gap between VLMs and real-world AR applications, our work also fosters more intelligent and interactive training experiences, benefiting industries that rely on precise, step-by-step assembly guidance.
B.3 Future Work
A key area for future research is improving VLMs for fine-grained spatial reasoning. Current models struggle with detailed object and state recognition, limiting their ability to provide accurate real-time feedback in AR training. Advancements in pretraining strategies, multimodal fusion techniques, and architectural design could enhance spatial understanding in AR-based assembly tasks.
Expanding the dataset is another critical direction. While our dataset focuses on LEGO-based assembly tasks as a simplified simulation, future work should explore real-world industrial assembly scenarios, such as furniture construction, electronic device assembly, and mechanical repairs. This would improve model generalization and practical applicability in professional training settings.
Additionally, integrating real-time multimodal interaction is essential for developing dynamic AR assistants. Future research should explore VLMs that can process live user input, track assembly progress, and provide adaptive feedback based on visual and textual cues. This would move beyond static dataset-based evaluation toward fully interactive AR guidance systems.
Finally, trust and explainability in human-AI collaboration remain crucial. Understanding how users interact with AI-powered AR assistants can help improve trust, usability, and transparency. Future work should investigate methods for making VLMs more interpretable and reliable, ensuring that AI-generated instructions align with human expectations and real-world assembly requirements.
Appendix C Implementation Details
C.1 Implementational setup
During data synthesis, we utilized MiniGPT4-v2777https://github.com/Vision-CAIR/MiniGPT-4 as the VLM, employing straightforward rules and human correction to create vision-language data. To evaluate the performance of benchmark models, we randomly sampled 25% of the dataset to serve as the test set. All inference and fine-tuning experiments were conducted on NVIDIA A100 SXM4 40GB GPUs. We set the learning rate to 1e-5 and the batch size to 16 for our experiments. The AdamW optimizer (Loshchilov & Hutter, 2017) was employed to facilitate efficient training, which spanned 3 epochs.
C.2 Query Templates
In this paper, we investigate three fine-grained vision-language tasks and have devised specific query templates for each, as demonstrated in Table 7.
T1: Scene Understanding
Token: [grounding]
Selection Rule: <step_symbol> with <step>. Instruction from this step.
Format-Directive: Tell me what to do for the current situation.
T2: Object Detection
Token: [object]
Selection Rule: <step_symbol> with <eop>, Word starts with “find” or “collect”. Instruction from this step.
Format-Directive: Providing the positions in the format: <p>object</p> {<|Xleft><Ytop><Xright><Ybottom>}, with X and Y coordinates normalized to [0,100]. <Xleft> and <Ytop> for the top-left corner. <Xright> and <Ybottom> for the bottom-right corner.
T3: State Detection
Token: [state]
Selection Rule: <step_symbol> with <step>. Previous <step_symbol> is <eop>. Instruction from this step and the previous step.
Format-Directive: Just tell me Yes or No.
Table 7: Query Templates for Fine-Grained VLM Tasks.
C.3 Dataset Details
C.3.1 Data Schema in JSON Format
This section presents the JSON data schema for integrating multimodal data (visual and textual) and fine-tuning VLM. The schema defines key components that structure and organize data efficiently, ensuring seamless preprocessing before final compilation.
  "instruction_id": 95,
  "text": [ "Find 2 transparent red flat tiles 1x2." ],
  "VLM": {
    "img_path": "...",
    "parts_img_path": "None",
    "id_sign": "Car",
    "step_num": 28,
    "step_class": "eop",
    "other_sign": "AG",
    "background_rect": "None",
    "color_list": [],
    "bound_color": [],
    "task_label": "object",
    "query": "Find 2 transparent red flat tiles 1x2.",
    "MiniGPTv2_output": "<p>2 transparent red flat tiles</p> {<28><17><73><50>}"
  "img": "...",
  "entities": [
    { "this_line": "Find 2 transparent red flat tiles 1x2." },
    { "verb": "Find" },
    { "det": "2" },
    { "colorinfo": "transparent red" },
    { "theme": "flat tiles 1x2" },
    { "suffix": "." } ]
Table 8: An example of a JSON schema for integrated multimodal data.
The key fields related to VLM in the JSON schema include:
  • img_path: Stores the local path of images downloaded via web scraping, pointing to the original files fetched from the web.
  • id_sign: Extracted from the image filename, this field identifies the sub-assembly session (e.g., “CAR” in this example).
  • step_num: Represents the sequence number of the assembly step (e.g., step 28), tracking the assembly’s progression.
  • step_class: A classification tag (e.g., “eop” for “end of part”) indicating an object-finding step in the assembly process.
  • MiniGPTv2_output: Stores processed output from MiniGPTv2, containing annotations or modifications essential for subsequent dataset creation.
  • parts_img_path, bound_color, color_list: Reserved for assembly steps classified as “step” and crucial for state detection dataset generation:
    • color_list: Records all possible boundary colors used throughout the assembly session.
    • bound_color: Consolidates the most probable boundary colors through a voting mechanism, ensuring consistency across the dataset.
To support fine-tuning, our default JSON format follows the Llava model’s fine-tuning data structure, ensuring an optimized and adaptable framework for general use cases. An example of the JSON schema for VLM fine-tuning is illustrated in Table 9.
  "id": "lego-60263-ocean-mini-submarine-readscr-12",
  "image": "sm01-island-0_step_Step%20%23%23E.png",
  "conversations": [
      "from": "human",
      "value": "[grounding] After I find 1 reddish brown chest bottom 2x4.
      Tell me what to do for the current situation."
      "from": "gpt",
      "value": "Put it horizontally on the table."
Table 9: An example of a JSON schema for VLM fine-tuning.
The full dataset, comprising JSON and PNG files, occupies approximately 11.8 GB. For specialized requirements, such as those for the Otter model, we convert all text and image data into Parquet files.
C.3.2 An Example of LEGO Instruction Manual
Figure 5 shows an example of a part of the instruction manual we crawled from the official LEGO website.
Figure 5: An example of LEGO instruction manual. It consists of a summary section at the beginning, followed by three sequential instruction steps. Each step includes textual instructions paired with corresponding images to guide the assembly process.
The LEGO audio instruction manual is a structured guide designed to assist blind and visually impaired users in assembling LEGO sets accurately. The key sections of a LEGO instruction manual include: (1) a summary providing general information, such as topics and candidate parts for assembly, followed by (2) a sequence of multimodal step instructions. Each step consists of textual instructions paired with an illustrative image, with key functional phrases, such as theme entities, highlighted for clarity provided by LEGO. With an average of 215.3 steps per manual, we segmented each manual into multiple assembly sessions, each consisting of a paired sequence: an object-seeking step followed by an object-assembly step.
C.3.3 Illustration of Data Creation.
Figure 6 illustrates the process of creating the dataset for task T2. It begins by iterating through object-seeking steps, where users locate specific objects. Next, images are processed to generate the object detection dataset, consisting of query-image-text triplets. Queries are formed using a template with the task token "[object]" and the current instruction. Responses include positional text indicating object coordinates, initially inferred using MiniGPT-v2. A composite image is then created by merging the target image with three randomly sampled images, and the coordinates are adjusted accordingly.
Figure 6: An illustration of object detection data creation.
Figure 7 illustrates the process of generating the dataset for task T3. The procedure begins by iterating through object-assembly steps, pairing each with its preceding step where users identified parts to assemble. Next, part segmentation is performed by detecting boundary colors using K-means clustering and Hue filtering, followed by manual correction. Boundaries are highlighted in distinct colors to mark assembly objects. Bounding boxes are then added to segmented objects, and natural perturbations are introduced by shifting assembled parts within the background box. Each correct assembly state serves as a positive sample, while three perturbed states act as negative samples. Finally, the dataset is constructed, where queries integrate task tokens, current and previous instructions, and responses indicate correct or incorrect assembly states.
Figure 7: An illustration of synthesized negative samples for state detection.
C.3.4 Examples in Data Quality Assessment
Figure 8 and 9 illustrate examples of the quality assessment for generated data in object detection (T2) and state detection (T3), respectively. We randomly selected 100 data examples per task to ensure a fair assessment. Three annotators were provided with comprehensive guidelines on how to evaluate each dataset based on multiple criteria, accompanied by relevant examples. The first five data points mentioned in the guidelines were used as a sanity check to ensure that the annotators correctly understood and applied the evaluation criteria. If annotators fail the sanity check, they undergo retraining until they pass. We evaluate annotations using the average scores of four criteria (i.e., entity disambiguity, boundary precision, state relevance, and state identifiability) and calculate Kappa values to assess annotation agreement.
Figure 8: An example of evaluating data quality of task T2.
Figure 9: An example of evaluating data quality of task T2.

### v2

@@ Introduction
GPT-4o: Ensure the transparent mailbox front 2x2 is securely attached by inserting the small shaft into the round brick 1x1 before proceeding with the next step.
Table 3: An example for understanding VLM challenges.
This suggests that annotators reached a moderate agreement on the clarity of the evaluated objects in relation to the corresponding scenes. The average state relevance and identifiability scores are 0.93 and 0.74 out of 1.00, respectively, with the Kappa scores 0.81 and 0.59. This indicates that annotators reached near-perfect agreement on the relevance of the generated states to the original while showing moderate agreement on the recognizable variation of those states. Besides, the first five data points mentioned in the guidelines were used as a sanity check to ensure that the annotators correctly understood and applied the evaluation criteria. This ensures the overall quality of the generated states.
Case Study of Fine-grained VLM Challenges
@@
Table 3 present an example showcasing the intuitive results of the prevailing VLMs on scene understanding. Compared with the reference, existing VLM models face key challenges: (1) Vague and generic assembly instructions.mPLUG-OWL2 and BLIP2 produce broad, non-specific placement directives (e.g., “Place the mailbox front 2x2 next to the block on the table” or “Put the mailbox front 2x2 horizontally on the table, clasp to the back”), falling short of the precise, step-by-step guidance found in the reference. Otter’s response is even more superficial, merely stating “build structure” without actionable detail. (2) Struggles with fine-grained entity recognition and attribute grounding.While models like Qwen-VL and MiniGPT-v2 correctly mention the “transparent mailbox front 2x2,” others, such as InstructBLIP, overlook critical attributes (e.g., referring only to “it” or “the mailbox front”). LlaVA, despite its verbosity, hallucinates details (e.g., “red liquid inside” and “blue sky”) absent from the scene, highlighting confusion in entity identification and property association. (3) Tendency to hallucinate or misinterpret assembly steps.MiniGPT-v2 and GPT-4o introduce extraneous or incorrect actions, such as “aligning with the short side of the plate” or “inserting the small shaft into the round brick 1x1,” which are not part of the actual assembly process. This reveals a shallow grasp of both scene context and the logical flow of assembly operations. Overall, while some models can partially identify the correct entities and occasionally mention relevant properties, none consistently generate instructions that are both factually accurate and contextually appropriate for fine-grained assembly tasks.
Related Work
Vision-Language Datasets for Assembly Tasks
@@
Various datasets have been developed to enhance assembly tasks by supporting key capabilities T1, T2, and T3. A comparative summary is presented in Table 4.
      Dataset       T1       T2       T3       Size
      COIN       ✓       ✗       ✓       11,827
      HowTo100M       ✓       ✗       ✓       23,611
      TEACh       ✓       ✗       ✓       3,215
      MIMIC-IT       ✓       ✗       ✗       2.8M
      HoloAssist       ✓       ✗       ✓       350
      EPIC-KITCHENS       ✗       ✓       ✓       89,977
      Assembly101       ✓       ✗       ✓       4,321
      Cross-task       ✗       ✓       ✓       4,713
      RareAct       ✗       ✓       ✓       7,607
      LEGO-VLM (Ours)       ✓       ✓       ✓       35,612
Table 4: Compariable datasets for assembly tasks towards AR training regarding the cabablities of T1, T2, T3. The symbols ✓and ✗indicate the presence or absence of each capability.
@@
COIN provides a strong foundation for sequential task analysis with comprehensive annotations, aiding research in multimodal learning (Tang et al. 2019). HoloAssist captures egocentric human-AI interactions using mixed-reality headsets, offering valuable real-world insights (Wang et al. 2023b). HowTo100M, with its extensive collection of 136 million video clips and transcribed narrations, is highly effective for text-to-video retrieval but lacks fine-grained object annotations (Miech et al. 2019). TEACh focuses on interactive dialogues in domestic environments, improving dialogue modeling but not prioritizing detailed object detection (Padmakumar et al. 2021).
@@
While some datasets excel in action recognition, they often lack the procedural depth necessary for effective training. Assembly101, with over 4,000 toy assembly videos, does not include step-by-step instructional details (Sener et al. 2022). CrossTask facilitates weakly supervised learning by leveraging narrations and step lists but lacks temporal annotations to clarify action sequences (Zhukov et al. 2019). EPIC-KITCHENS provides extensive annotations on unscripted kitchen activities, yet it does not offer structured procedural guidance (Damen et al. 2018). RareAct presents unique interactions, challenging models to interpret complex actions without explicit instructions (Miech et al. 2020).
@@
Stanescu et al. (2023a) introduce a state-aware prior that significantly improves object detection in assembly tasks like furniture and Lego construction. However, it lacks textual descriptions of object states and does not address fine-grained detection (Stanescu et al. 2023b). Meanwhile, MIMIC-IT contributes 2.8 million multimodal instruction-response pairs, enriching conversational modeling but lacking scenario-specific descriptions that would enhance task guidance (Li et al. 2023a).
@@
Towards Fine-grained Vision-Lanuage Models
@@
Recent advancements in vision-language models (VLMs) have broadened their applications, enhancing accessibility, search indexing, and interactive content retrieval (Fan et al. 2024; Zhang et al. 2024; Du et al. 2022; Gan et al. 2022). These models generate captions, describe images (Zhou et al. 2020; Hu et al. 2022), facilitate visual QA (Bazi et al. 2023). For instance, they can accurately identify objects in images based on phrases like “the red car” (Subramanian et al. 2022). Furthermore, these models facilitate conversational interactions about visual elements (Chen et al. 2022), which makes virtual assistants more adept at handling inquiries related to images and videos, which makes virtual assistants more adept at handling inquiries related to images and videos. Additionally, VLMs can support multilingual content (Gwinnup and Duh 2023), benefiting assistive technologies and education (Chi et al. 2020; Wang et al. 2024a).
@@
Despite progress, VLMs often lack fine-grained modeling for scene understanding, object recognition, and error detection. While CLIP demonstrates strong generalization (Radford et al. 2021), models like EfficientVLM (Wang et al. 2023a), MiniGPT-v2 (Chen et al. 2023), and Qwen-VL (Bai et al. 2023) struggle with intricate tasks. OSCAR (Li et al. 2020) and VisionLLM (Wang et al. 2024b) improve image-text alignment but face challenges in context-specific adaptation. The Otter model advances sequential task management yet falls short in detailed analysis (Li et al. 2023b).
@@
Efforts in procedural video representation (Zhong et al. 2023) and instructional task graphs (Ashutosh et al. 2024) highlight potential VLM integration but reveal gaps in error correction. To address these challenges, we propose a vision-language architecture optimized for fine-grained tasks, advancing training and benchmarking of VLMs.
Discussion
Implications and limitations. This work addresses a critical gap in the development of intelligent, instruction-following multimodal assistants by focusing on fine-grained procedural understanding—an ability crucial for equitable access to education and skill development in a technologically evolving society. While our benchmark is partially synthetic and does not yet support 3D scene understanding, it represents an important step toward scalable, accessible AI systems that can interpret complex visual-textual instructions. The implications span beyond research: such systems could democratize hands-on learning by enabling visually impaired users to access assembly tasks through multimodal feedback, and could revolutionize technical education and vocational training by offering consistent, language-grounded support. Our findings further highlight the current limitations of VLMs in precise reasoning and verification, pointing toward future directions in responsible AI development that benefits society at large—including learners, workers, and underserved communities.
Ethical Considerations. We recognize the ethical implications of developing VLMs for user interactions, so addressing these concerns is essential. To ensure ethical standards, we use open-source VLMs as benchmarks and prioritize user-centric design such as inclusiveness for diverse users.
Conclusion
@@
Vision-language models (VLMs) have significantly advanced in recent years, enabling AI systems to interpret and generate textual descriptions of visual content. However, despite their success in general vision-language tasks, these models often struggle with fine-grained understanding, particularly in structured instructional scenarios. This research explores fine-grained vision-language modeling for manual-guided LEGO assembly tasks, focusing on scene understanding, object detection, and state detection—areas where VLMs often struggle. We developed a specialized dataset from LEGO instruction manuals, designing fine-grained tasks to evaluate VLM performance in tracking assembly sequences and interpreting instructions. Through a one-for-all architecture, we assessed existing models and found them lacking in precision for instructional tasks, highlighting the need for improvement. Future work will analyze model failures and develop enhanced VLMs with better accuracy and contextual understanding. Beyond technical advancements, our research aims to empower blind and visually impaired individuals by enabling AI-driven learning tools, promoting greater accessibility and independence.